-6x=-5y+8
2x+6y=28

Answer by cleomenius(959)

I assume we are solving simultaneous equations.
-6x = -5y + 8
2x + 6y = 28
add 5y to each side in equation 1.
multiply 2nd equation by 3.
-6x + 5y = 8
6x + 18y = 84
add equations
23y = 92.
y = 4.
solve for x in equation 1.
-6x = -5(4) + 8
x = 2
substitute x and y into equation 2.
2(2) + 6(4) = 28.
This does check.
Cleomenius.